OpenStudy (anonymous):

find an expression for a rational function f(x) that satisfies the conditions: a slant asymptote of y = 2x, vertical asymptote at x = 1, and contains the point (0,6)

OpenStudy (anonymous):

@wio can you help?

OpenStudy (anonymous):

I suppose I can try.

OpenStudy (anonymous):

To get vertical asymptote you want something like \[ \frac 1{x-1} \]

OpenStudy (anonymous):

For the slant asymptote you want to add \(2x\) to it: \[ 2x+\frac 1{x-1} \]

OpenStudy (anonymous):

Now that numerator can be what we want it to me... let me change: \[ 2x+\frac {a}{x-1} \]

OpenStudy (anonymous):

We plug in the point to solve for \(a\). \[ 2(0)+\frac{a}{(0)-1}=6 \]

OpenStudy (anonymous):

Solving for \(a\) gives us: \[ \frac{a}{-1}=6\implies a=-6 \]

OpenStudy (anonymous):

So now that we have our equation, we simplify: \[ f(x) = 2x-\frac{6}{x-1}=\frac{2x(x-1)-6}{x-1}=\frac{2x^2-2x-6}{x-1} \]

OpenStudy (anonymous):

@studygeek15 It's basically all about reverse engineering.

OpenStudy (anonymous):

Thank you so much!